# Phase locking and multistability in the topological Kuramoto model on cell complexes

### v1

@@ Acknowledgements.
D.W. and I.B. gratefully acknowledge support by the German Federal Ministry of Research, Technology and Space (Bundesministerium für Forschung, Technologie und Raumfahrt) via the grant number 03SF0751. M.T.S acknowledges support by the European Union (ERC, HIGH-HOPeS, 101039827). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Eigenvalues of the Jacobian
${\boldsymbol{J}}{\boldsymbol{v}}=\lambda{\boldsymbol{v}}$
$\Rightarrow\;$ $\lambda=\frac{{\boldsymbol{v}}^{\top}{\boldsymbol{J}}{\boldsymbol{v}}}{{\boldsymbol{v}}^{\top}{\boldsymbol{v}}}\,.$
@@
If the stationary state satisfies the conditions in Eq. (11), this implies
$\lambda<0.$

### v2

@@ Acknowledgements.
We thank Ginestra Bianconi for stimulating discussion. D.W. and I.B. gratefully acknowledge support by the German Federal Ministry of Research, Technology and Space (Bundesministerium für Forschung, Technologie und Raumfahrt) via the grant number 03SF0751. M.T.S acknowledges support by the European Union (ERC, HIGH-HOPeS, 101039827). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Eigenvalues of the Jacobian
${\boldsymbol{J}}{\boldsymbol{v}}=\lambda{\boldsymbol{v}}\qquad\Rightarrow\qquad\lambda=\frac{{\boldsymbol{v}}^{\top}{\boldsymbol{J}}{\boldsymbol{v}}}{{\boldsymbol{v}}^{\top}{\boldsymbol{v}}}\,.$
@@
If the stationary state satisfies the conditions in Eq. (11), this implies $\lambda<0$.
Appendix B Proof of Eq. (23)
To proof Eq. (23), we rewrite Eq. (22) in components and use that $|f_{\pm}|\leq\pi/2$ and obtain
[…]
We recall that the matrices $\boldsymbol{C}$} are closely related to the boundary matrices. We can choose a basis of the kernel such that the $k$th column of ${\boldsymbol{C}}_{[n+1]}$ (${\boldsymbol{C}}_{[n]}^{\top}$) is equal to the $k$th column of ${\boldsymbol{B}}_{[n+2]}$ (${\boldsymbol{B}}_{[n-1]}^{\top}$) unless $k$ corresponds to the harmonic component. A further simplification is possible by noting that the entries of the boundary matrix have the values 1,0,+1-1,0,+1. We then define $m_{k}^{[+]}$ as the number of non-zero elements of the $k$th column of ${\boldsymbol{B}}_{[n+2]}$ and $m_{k}^{[-]}$ as the number of non-zero elements of the $k$th row of ${\boldsymbol{B}}_{[n-1]}$. We then obtain the bounds given in Eq. (23)for every $k$ that does not corresponds to the harmonic component.
Notably, the number of non-zero elements can be related to the number of adjacent cells and thus enables a geometric interpretation. The number $m_{k}^{[+]}$ counts the number of cells of dimension $n+1$ in the boundary of the $k$th cell of dimension $n+2$. The winding number $z_{k}^{[+]}$ counts how often the phases $\boldsymbol{\theta}$$^{[+]}$ wind by $2\pi$ when following all cells in this boundary. Likewise, the number $m_{k}^{[-]}$ counts the number of cells of dimension $n$ that have the $k$th cell of dimension n1n-1 in its boundary.
Appendix C Proof of equations (28)
We prove the equations (28) using a recurrence relation that is derived from the fact that $d$-simplex is contractible and the rank-nullity theorem. A $d$-simplex is contractible, which means that all homology groups are trivial except $H_{0}\cong\mathbb{R}$,
$H_{n}=0,\quad n>0,$ (29)
where a homology group is the quotient vector space
$H_{n}=\frac{\ker{\boldsymbol{B}}_{[n]}}{\text{im }{\boldsymbol{B}}_{[n+1]}}$ (30)
implying
$\dim H_{n}=\dim\ker{\boldsymbol{B}}_{[n]}-\dim\text{im }{\boldsymbol{B}}_{[n+1]}.$ (31)
Since ${\boldsymbol{B}}_{[n]}{\boldsymbol{B}}_{[n+1]}=0$, every element in ${\boldsymbol{B}}_{[n+1]}$ lies in $\ker{\boldsymbol{B}}_{[n]}$,
$\text{im }{\boldsymbol{B}}_{[n+1]}\subseteq\ker{\boldsymbol{B}}_{[n]}$ (32)
For $n>0$, since $H_{n}=0$, we have
$\dim\ker{\boldsymbol{B}}_{[n]}=\dim\text{im }{\boldsymbol{B}}_{[n+1]}$ (33)
Combining Eqs. (32) and (33), we find for every $n>0$
$\text{im }{\boldsymbol{B}}_{[n+1]}=\ker{\boldsymbol{B}}_{[n]}.$ (34)
We now apply the rank nullity theorem to ${\boldsymbol{B}}_{[n]}:\mathbb{R}^{N_{[n]}}\to\mathbb{R}^{N_{[n-1]}}$ which yields
$\dim\ker{\boldsymbol{B}}_{[n]}=N_{[n]}-\text{rank }{\boldsymbol{B}}_{[n]}.$ (35)
We thus obtain a recurrence relation valid for $n>0$
$N_{[n]}=\text{rank }{\boldsymbol{B}}_{[n+1]}+\text{rank }{\boldsymbol{B}}_{[n]}=\binom{d+1}{n+1}\,$, (36)
where we used the fact that for a $d$-simplex, the number of $n$-dimensional faces, formed by choosing $n+1$ vertices from $d+1$ available ones, is $N_{[n]}=\binom{d+1}{n+1}$. Applying the recurrence relation for $n=d,\ldots,1$ together with Pascal’s identity then yields Eq. (28). As a last step, for $n=0$, we have to slightly adapt the recurrence relation because
$\dim H_{0}=1=\dim\ker{\boldsymbol{B}}_{[0]}-\dim\text{im }{\boldsymbol{B}}_{[1]}.$ (37)
Nevertheless, this yields the same results as Eq. (28) for $\boldsymbol{z}$$^{[+]}$. Finally, there is no $\boldsymbol{z}$$^{[-]}$ for $n=0$ which concludes the proof.